OpenStudy (anonymous):

Write the following as a single power of 2? 2^1-x/4^x+2

OpenStudy (zehanz):

If you write it in the equation editor:\[\frac{ 2^{1-x} }{ 4^{x-2} }\]This is already more clear. Now remember that 4=2², so write it as:\[\frac{ 2^{1-x} }{ (2^2)^{x-2} }\] Use these rules:\[(a^b)^c=a^{bc}\]and\[\frac{ a^b }{ a^c }=a^{b-c}\]to finish it off!